An inventor claims to have devised a closed cyclic engine which exchanges heat with cold and hot reservoirs at 25 and 350 ◦C, respectively, and produces 0.45 kJ of work for every kJ of heat extracted from the hot reservoir. Is this claim believable?
Answer #1

As given data,

Tc = 25 0C And Th = 350 0C

Work Produced per KJ of Heat Extracted from Hot Reservoir = 0.45KJ = Efficiency

If the device is Carnot cycle then efficiency will maximum and value will = 1 - (Tc/Th )

=1- (25/350) = 0.928

Hence, this type of device is possible and the claim is also belevable.